Ralston Manufacturing began operations in January 2018. Ralston sells computers that carry a two-year manufacturers warranty against defects in workmanship. Ralstons management projects that 3% of the computers will require repair during the first year of the warranty while approximately 6% will require repair during the second year of the warranty. The computers sell for $1,100 each. The average cost to repair a computer is $150. The company sells the computers to retail customers who must pay a 6% sales tax. Sales and warranty information for 2018 and 2019 are as follows: 3. 2018: Sold 500 computers on account; incurred warranty expenditures of $2,300. 2019: Sold 800 cameras on account, incurred warranty expenditures of $8,000. INSTRUCTIONS a. b. Prepare journal entries that summarize the sales and all aspects of the warranty for 2018. Prepare journal entries that summarize the sales and all aspects of the warranty for 2019.
Answer #1

Sales for2018 =500×1100=550000

Sales tax = 500×1100×6%=33000

Estimated warranty expense=550000×3%=16500

For 2019 for sales made in 2018 =500×6%×1100=33000 according to matching concept it should be recognised in 2018(year of sale)

For 2019

Sales=800×1100=880000

Sales tax= 880000×6%=52800

Estimated warranty expense = = 880000×3%=26400+ 880000×6%=52800 to be incurred in second year

Journal entries for year 2018 and 2019 reflecting above trasactrans

Date particulars debit credit
2018

customers account debit

To sales

To sales tax payable

Being sales made on account and sales tax collected

583000

550000

33000

2018

warranty expense account debit

To estimated warranty liability

Beign estimated warranty recognised

49500

49500

2018

estimated warrenty liability debit

To inventory

Beign warranty claimed by customers

2300

2300

2019

customers account debit

To sales account

To sales tax payable

Beign sales made on account and sales tax collected

932800

880000

52800

2019

warrenty expense account debit

To estimated warranty liability

Beign estimated warranty recognised

79200

79200

2019

warranty liability account debit

To inventory

Beign claims received from customers

8000

8000